When the price of good X falls from . 10 to . 9, the demand for good Y
Increase from 20 Kg. to 25 Kg.
a) What is the cross elasticity of demand of good Y for good X?
b) Are goods X and Y compliments or substitutes?
1
Expert's answer
2011-12-29T08:10:29-0500
Ed = (10-9)/(20-25)*45/19 = 1/(-5)*45/19 = -9/19
Two goods that complement each other show a negative cross elasticity of demand: as
the price of good Y falls, the demand for good X rises.
So the goods are complements.